OpenStudy (anonymous):

11n-4(2n-3)=18+5(n+2)

OpenStudy (anonymous):

distribute the -4 through the parenthesis and distribute the 5 in through the parenthesis. 11n - 4(2n - 3) = 18 + 5(n + 2) 11n - 8n + 12 = 18 + 5n + 10 now combine like terms....get the x's on one side and everything else on the other side. Then, just solve for x. Can you do that ?

OpenStudy (anonymous):

3n+12=5n+28

OpenStudy (anonymous):

This is where I get stuck. Do I subtract -12 or -28 from both sides?

OpenStudy (anonymous):

either one....you can subtract 12 from both sides or you can subtract 28 from both sides. Your main goal is to get the x's on one side, it doesn't matter which side. Either way you do it, whether it is subtracting 12 or subtracting 28, you should get the same answer.

OpenStudy (anonymous):

Ohhhh OK Thanks, helps if the teachers tells you that bit of info

OpenStudy (anonymous):

try to do the problem and I will check it

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

11n - 4(2n - 3) = 18 + 5(n + 2) 11n - 8n + 12 = 18 + 5n + 10 3n + 12 = 5n + 28 - 28= -28 3n -16 = 5n -3n =-3n -16 =-2n -2 =-2 8 = n

OpenStudy (anonymous):

Kelliegirl33 are you still here? The screen says your replying but its been 10minutes

OpenStudy (anonymous):

11n - 8n + 12 = 18 + 5n + 10 (combine like terms) 3n + 12 = 28 + 5n 3n + 12 - 28 = 28 + 5n - 28 3n - 16 = 5n 3n - 3n - 16 = 5n - 3n -16 = 2n -16/2 = n -8 = n or... 3n + 12 = 28 + 5n 3n + 12 - 12 = 28 + 5n - 12 3n = 16 + 5n 3n - 5n = 16 + 5n - 5n -2n = 16 n = 16/-2 n = -8

OpenStudy (anonymous):

do you see where your mistake was ?

OpenStudy (anonymous):

That's what I'm trying to figure out. I got a positive 8

OpenStudy (anonymous):

OIC I used a -2

OpenStudy (anonymous):

Lets start here.... 3n - 16 = 5n....you subtracted 3n... -3n - 3n ------------- -16 = 2n (5n - 3n = 2n not -2n) next, you have to divide by 2 to get n by itself -16/2 = 2n/2 -16/2 = n -8 = n do you understand now ?

OpenStudy (anonymous):

as you can see, you can either subtract the 12 or subtract the 28, and you will get the same answer. Just get the x's on one side, either side, and solve for x.

OpenStudy (anonymous):

Thank you so much for your help!

OpenStudy (anonymous):

your very welcome....that is what I am here for :) you got any more questions ?

OpenStudy (anonymous):

Not at the moment. Still have about 10 more math problems to do until Im done.

OpenStudy (anonymous):

you can do them....your smart. Oh....by the way....you got a cool user name. I guess I could have made my user name softball junky, but I chose kelliegirl instead.

OpenStudy (anonymous):

LOL! Yes Love me some Soccer! :-)

OpenStudy (anonymous):

thats cool.....anymore problem, just post them...people are here to help :)

OpenStudy (anonymous):

Thanks!